if 34.0356 is expressed in the form p/q where p and q are co prime integers then what can you say about the factorisation of q?
Answer by stanbon(57361)

if 34.0356 is expressed in the form p/q where p and q are co prime integers then what can you say about the factorisation of q?
----
= 340356/10^4
---
= 170158/5000
---
= 85079/2500
==================
Cheers,
Stan H.